How do I work this Problem out 2(3h-2)-2h=2(h-3)h=?
Thanks
: How do I work this Problem out 2(3h-2)-2h=2(h-3)h=?
Thanks

Answer by benni1013(141)

First breathe this looks like a hard problem but isn't.
2(3h-2)-2h=2(h-3)
Use the distributive property to obtain the following.
6h-4-2h=2h-6
Combine like terms
4h-4=2h-6
Anytime when you do a long equation you will always end up with the above format. Take the lower coefficent and get it to the other side.
2h-4=-6
Add 4
2h=-2
Divide by 2
h=-1